# Dynamical correlations and nonequilibrium sum rules in photodoped Hubbard ladders

### v2

@@ III.2 Equal time correlations
Spin-spin correlations are strongly suppressed by the pump, with the extent of the suppression and the details of the relaxation afterwards depending on both the energy imparted and the pump direction. In Fig. 3 we plot the staggered spin-spin correlations along one leg $(-1)^{x}\langle S^{z}_{i,l}S^{z}_{i+x,l}\rangle$ as a function of separation $x$ and time $t$, for three different pump protocols.
The ground state of the $U=8$ half filled Hubbard ladder is a Mott insulator with strong antiferromagnetic correlations and a spin-spin correlation length $\xi_{s}\approx 4.2$ (in agreement with Ref. [27]). This is reflected at $t=0$ in each panel of Fig. 3, where the spin-spin correlations extend to large distances using a logarithmic colour scheme. Note that the results for $x=0$ correspond to those in Fig. 2, but the suppression of onsite magnetization is less visible on a log scale. All three panels of Fig. 3 show suppression of spatial correlations during the pump. This is effect is fastest for the leg direction, $A_{0}=1$ pump (Fig. 3 upper panel), where correlations beyond nearest neighbour are suppressed by an order of magnitude by the midpoint of the pump. The plots for the less energetic, $A_{0}=0.45$ leg direction and $A_{0}=1$ rung direction, pumps in the middle and lower panels of Fig. 3 show suppression of longer-range spin correlations that continue past the point of peak pump amplitude. All three panels display a limited recovery of spin correlations sometime after the pump has peaked, with correlations growing outwards from $x=0$. The recovered spin-spin correlations are significantly weaker than in equilibrium, an effect that is seen clearly at next-nearest neighbour and beyond, $|x|\geq 2$. In addition, for the $A_{0}=0.45$ leg direction and $A_{0}=1$ rung direction pumps there is a local minimum in the spin correlations at $x=2$ in the post pump region.
Figure 3: Staggered spin-spin correlations $(-1)^{x}\langle S^{z}_{i,l}S^{z}_{i+x,l}\rangle$ along a leg as a function of separation $x$ and time $t$. The white dashed line indicates the time at which the pump is turned off. Upper panel: leg direction pump with $A_{0}=1$. Middle panel: leg direction pump with $A_{0}=0.45$. Lower panel: rung direction pump with $A_{0}=1$.
Figure 4: Connected charge (density-density) correlations $|\langle n_{i,l}n_{i+x,l}\rangle_{c}|$ along a leg as a function of separation $x$ and time $t$. The white dashed line indicates the time at which the pump is turned off. Upper panel: leg direction pump with $A_{0}=1$. Middle panel: leg direction pump with $A_{0}=0.45$. Lower panel: rung direction pump with $A_{0}=1$.
@@
The charge correlation length in the ground state of the half filled Hubbard ladder with $U=8$ is much shorter than the spin correlation length, and this is evident in Fig. 4 at $t=0$. In contrast to the spin correlations, the spatial charge correlations only grow with time during the pump. This growth is strongest for the $A_{0}=1$ leg direction pump (top panel), and weakest for the pump along the rung direction. The growth proceeds without any apparent sharp features associated with the time at which the pump amplitude peaks or the time at which the pump ends. Unlike the spin correlations, the charge correlations within the ‘light cone’ appear to decrease monotonically for small $x$, and there is no distinct minimum at $x=2$ in the charge correlations for the less energetic pump protocols. Hence, there are distinct qualitative differences between the spatial spin and charge correlations, despite the simple relation between the local magnetisation and doublon number.
@@
Table 1: Correlation lengths for spin, $\xi_{s}$, and charge, $\xi_{c}$, in equilibrium and at $t=10$ after pumping. We fit the correlation length between $x=6$ and $x=12$ using the data shown in Fig. 5. We include 2 significant figures, fit uncertainties are $<10^{-2}$.
@@
The spin correlation length $\xi_{s}$ is reduced by approximately one lattice unit after the pump. However, the most notable difference between the equilibrium and nonequilibrium spin correlations is the very fast decay of the latter at short range, before the exponential decay characterised by $\xi_{s}$ takes over. This behaviour is evident for $x\leq 3$ in the upper panel of Fig. 5 where, although the spin correlations at $x=0$ (i.e. the squared magnetizations) are of similar magnitude for the equilibrium and nonequilibrium data, by $x=3$ there is an order of magnitude difference between the equilibrium and nonequilibrium cases. This effect is even more pronounced for the $A_{0}=0.45$ leg direction pump and the rung direction pump because of the local minimum at $x=2$. Thus, the value of $\xi_{s}$ alone does not capture the extreme weakness of correlations beyond nearest neighbour in the nonequilibrium state.
The inset of the lower panel of Fig. 5 and the data in Table 1 show that connected charge correlations after pumping are strongly enhanced relative to equilibrium. Although the post pump spin and charge correlation lengths are similar, their short range behaviours are different. For charge correlations there is no additional strong suppression on going from $x=1$ to $x=2$, and there is no local minimum at $x=2$. The increase in $\xi_{c}$ can be interpreted as a consequence of the conservation of charge: the pump increases the doublon occupancy of a site, which must also reduce the neighbouring charge density, and thus leads to an enhanced anticorrelation.
@@
In Fig. 6 we show the correlations at the antiferromagnetic wave vector $(\pi,\pi)$. We also include data for finite, $L=16$, rung ladders with open boundary conditions, using the two central rungs as a proxy for the two rung unit cell of approximate infinite system. The leg direction $A_{0}=1$ pump shows the fastest suppression of the antiferromagnetic peak, with $S^{zz}(\pi,\pi,t)$ essentially reaching its late time value by $t\approx 2$, during the pump. For the lower fluence, $A_{0}=0.45$, leg direction pump the suppression of $S^{zz}(\pi,\pi,t)$ is slower, reaching its late time value close to $t=\pi$ when the pump ends. The rung direction pump leads to the slowest suppression of the three protocols, with $S^{zz}(\pi,\pi,t)$ decreasing until $t\approx 4$ after the pump has ended. At $t=0$ the $L=16$ rung data differ from the $L=\infty$ value of $S^{zz}(\pi,\pi,t)$ due to finite size effects. As the spin correlation length decreases, the difference between a finite and infinite system becomes less distinct, and for all three pump protocols there is good agreement between the $L=16$ and $L=\infty$ data by time $t\approx 1$. This agreement continues to improve until $t\approx 5$, at which point small differences between the $L=16$ and $L=\infty$ curves become apparent again. This is consistent with the partial recovery of longer range staggered spin correlations seen in Fig 3 which increases the influence of the boundary in the finite system. We note that there is still very good agreement at $t=10$ between the $L=16$ and $L=\infty$ data for the three protocols. As $\xi_{c}\approx\xi_{s}$, these results suggest that an $L=16$ system is a good approximation for an infinite ladder during and after the pump.
@@ IV Nonequilibrium dynamical structure factor
Here the $g(t_{i},t)$ are windowing functions associated with the probe, which we assume to have a Gaussian form,
@@
where $\sigma_{X}$ is the probe width. $C(\mathbf{k},t_{1},t_{2})=\langle O_{\mathbf{k}}(t_{1})O^{\dagger}_{-\mathbf{k}}(t_{2})\rangle$ is a two time correlation function in momentum space. For our purposes $O_{\mathbf{k}}=S^{z}_{\mathbf{k}}$ or $n_{\mathbf{k}}$, respectively the spin $z$-component or charge density operator in momentum space.
@@
×$\times\left[C(\mathbf{k},t+\tau,t)+C(\mathbf{k},t,t-\tau)\right].$ (6)
This approximation is reasonable if $C(\mathbf{k},t+\tau,t)$ varies slowly with $t$ relative to the scale $\sigma_{X}$. In our results we set $\sigma_{X}=1$. We provide more details in Appendix B including a comparison that shows there is good agreement between the approximate form Eq. 6 and the full form Eq. 4, even during the pump when the conditions for the approximation are not strictly met.
@@ IV.1 Nonequilibrium sum rules
In this subsection we show that the combined total spectral weight of the nonequilibrium dynamical spin and charge responses is time-independent. Relations of this type, between dynamical quantities and a constant, are called sum rules. The total spectral weight in a channel is given by integrating Eq. 4 for $I(\mathbf{k},\omega,t)$ over all $\mathbf{k}$ and $\omega$,
@@ IV.2 Spin dynamical structure factor
Figure 7: Nonequilibrium dynamical spin structure factor at $k_{y}=\pi$ for a pump with $A_{0}=1$ in the leg ($\parallel$) direction. The leftmost column shows the response in equilibrium before the pump. The middle and rightmost columns show results during and after the pump respectively. The black open circles mark maxima in the response at a given wavevector $k_{x}$. The corresponding plots for the leg direction pump with $A_{0}=0.45$ and rung direction pump with $A_{0}=1$ are qualitatively very similar.
We compute the dynamical spin response $S(k_{x},k_{y},\omega,t)$ using Eq. 6 before pumping, mid way through the pump at $t=\pi/2$ and after the pump at $t=10$. Fig. 7 shows the spin response for $k_{y}=\pi$ as a function of $\omega$ and $k_{x}$ for a leg direction pump with $A_{0}=1$ at three different times: $t=-10$ before pumping (equivalent to the equilibrium result), during the pump at $t=\pi/2$, and at $t=10$ after pumping. We do not show the results for the $A_{0}=0.45,\parallel$ and $A_{0}=1,\perp$ pumps as they are qualitatively very similar when plotted in this way.
@@
In equilibrium, the spin excitations of the Hubbard ladder at half filling and large $U$ are well described by those of a Heisenberg ladder with antiferromagnetic exchange interactions. This two-leg spin ladder has a ground state characterised by singlet formation on rungs [26]. Spin excitations are formed by breaking rung singlets, leading to massive (gapped) dispersing triplets. There is a spin gap and no true long range order, and the system is sometimes described as a spin liquid. However, because the spin gap $\Delta_{s}\approx 0.12$ is small and the antiferromagnetic correlations are somewhat long ranged, the low energy dynamics are approximately those of a spin-1 magnon, with a sharply defined mode that has an almost linear dispersion relation as it approaches zero energy near $\mathbf{k}=(\pi,\pi)$ [35, 36]. This behaviour (convolved with the probe resolution) is seen in the left-hand panel of Fig. 7. The middle panel of the same figure shows that during the pump the spin response near $\mathbf{k}=(\pi,\pi)$ is reduced in intensity, in agreement with the real time results of the previous section. There is also a notable broadening of the response. At later times after pumping (right-hand panel of Fig. 7) the antiferromagnetic peak is substantially reduced in intensity, the response is broader for all $k_{x}$, and the dispersion has flattened.
@@
Figure 8: Nonequilibrium dynamical spin structure factor at $k_{y}=0$ for the $A_{0}=1$ leg direction pump, before, during, and after the pump. The black open circles mark maxima in the response at a given wavevector $k_{x}$. The corresponding plots for the leg direction pump with $A_{0}=0.45$ and rung direction pump with $A_{0}=1$ are qualitatively very similar.
@@
Fig. 8 shows the equivalent dynamical spin response for the $A_{0}=1$ leg direction pump along $k_{y}=0$ (as for $k_{y}=\pi$, the other two pump protocols give very similar results). The response along $k_{y}=0$ vanishes as $k_{x}$ approaches 00 (because $\sum_{i,l}S^{z}_{i,l}=0$ is a conserved quantity) and is strongest at $k_{x}=\pi$. After the pump the response at $(\pi,0)$ is weaker, although this suppression is less significant than at $(\pi,\pi)$. There is also someenhancement of the spectral weight close to $k_{x}=0$ after pumping, relative to the equilibrium case.
@@
Figure 9: Magnetic correlations with $\mathbf{k}=(\pi,\pi)$, at $t=10$ after pumping along the legs ()($\parallel$)( ∥ ) or rungs ()perpendicular-to($\perp$)( ⟂ ). The inset shows the strong antiferromagnetic peak in equilibrium.
@@
Figure 10: Magnetic correlations with $\mathbf{k}=(\pi,0)$, at $t=10$ after pumping along the legs ()($\parallel$)( ∥ ) or rungs ()perpendicular-to($\perp$)( ⟂ ).
@@
More precise information on changes in the energy profile of the spin response is provided by taking cuts at fixed wave vectors. In Figs. 9 and 10 we compare the equilibrium and $t=10$ post pump responses for $\mathbf{k}=(\pi,\pi)$ and $\mathbf{k}=(\pi,0)$ respectively  [^1] The very small amount of negative spectral weight visible for two of the curves in Figs. 9 and 10 is a consequence of the finite window used in the frequency Fourier transform. This negative weight could be removed by using a larger time window at a significant computational cost.. The equilibrium response at both wave vectors is essentially a resolution limited Gaussian peak. Other than the overall weakening of the response, the most notable result of the pump at $\mathbf{k}=(\pi,\pi)$ is that the broadening of the peak occurs in an asymmetric manner with a tail towards higher energies, in contrast to the symmetric broadening that is expected in a simple scattering lifetime approximation.
@@
Asymmetric broadening of a single particle mode in the dynamical spin response has been described previously in gapped spin systems, with triplet excitations and hard-core interactions, at low but finite temperatures [38, 15, 16, 17, 39]. At zero temperature the single particle mode in these systems is a result of interband transitions from the ground state, by breaking singlets and creating states with a single propagating triplet. As temperature increases, triplet states become thermally populated and due to strong interactions they dress the single particle mode in a nontrivial manner, leading to an interband response with an asymmetric lineshape. Moreover, the existence of thermally populated triplets reduces the joint density of states for interband transitions, so the spectral weight associated with the single mode response decreases. These effects are accompanied by the emergence of a new spin response at $\omega\sim 0$ due to intraband transitions (scattering of the probe from thermally occupied states). This increased spectral weight at low energies is known as a Villain mode [40], and it compensates for the decrease in the single mode weight. The total integrated spin response then remains a conserved quantity as expected for a pure spin model.
@@
The total spin spectral weight $W_{S^{z}}$ can be determined using Eq. 10 from the doublon number computed with iTEBD, or by integrating the dynamical spin response for the 16 rung system shown in Figs. 7 and  8 over all $\omega$ and $\mathbf{k}$. We find excellent agreement between both these methods, which provides further confirmation of the accuracy of the approximation Eq. 6 and the use of the 16 rung system. For the $A_{0}=1$ leg direction pump the ratio of the spin weight after the pump to that in equilibrium is $W_{S^{z}}(t=10)/W_{S^{z}}(t=-10)=0.76$. For the $A_{0}=1$ rung direction pump the corresponding figure is $W_{S^{z}}(t=10)/W_{S^{z}}(t=-10)=0.86$. Therefore, the integrated spin response decreases significantly after the optical pump, indicating a substantial transfer of spectral weight to the charge sector.

### v3

@@ III.2 Equal time correlations
Spin-spin correlations are strongly suppressed by the pump, with the extent of the suppression and the details of the relaxation afterwards depending on both the energy imparted and the pump direction. In Fig. 3 we plot the staggered spin-spin correlations along one leg $(-1)^{x}\langle S^{z}_{i,l}S^{z}_{i+x,l}\rangle$ of an infinite ladder, as a function of separation $x$ and time $t$, for three different pump protocols.
The ground state of the $U=8$ half filled Hubbard ladder is a Mott insulator with strong antiferromagnetic correlations and a spin-spin correlation length $\xi_{s}\approx 4.2$ (in agreement with Ref. [43]). This is reflected at $t=0$ in each panel of Fig. 3, where the spin-spin correlations extend to large distances (as depicted using a logarithmic colour scheme).
Note that the results for $x=0$ correspond to those in Fig. 2, but the suppression of onsite magnetization is less visible on a log scale. All three panels of Fig. 3 show suppression of spatial correlations during the pump. This is effect is fastest for the leg-direction, $A_{0}=1$ pump (Fig. 3 upper panel), where correlations beyond nearest neighbour are suppressed by an order of magnitude by the midpoint of the pump. The plots for the less energetic, $A_{0}=0.45$ leg-direction and $A_{0}=1$ rung-direction, pumps in the middle and lower panels of Fig. 3 show suppression of longer-range spin correlations that continue past the point of peak pump amplitude. All three panels display a limited recovery of spin correlations sometime after the pump has peaked, with correlations growing outwards from $x=0$. The recovered spin-spin correlations are significantly weaker than in equilibrium, an effect that is seen clearly at next-nearest neighbour and beyond, $|x|\geq 2$. In addition, for the $A_{0}=0.45$ leg-direction and $A_{0}=1$ rung-direction pumps there is a local minimum in the spin correlations at $x=2$ in the post pump region.
Figure 3: Staggered spin-spin correlations $(-1)^{x}\langle S^{z}_{i,l}S^{z}_{i+x,l}\rangle$ along a leg of an infinite ladder as a function of separation $x$ and time $t$. The white dashed line indicates the time at which the pump is turned off. Upper panel: leg-direction pump with $A_{0}=1$. Middle panel: leg-direction pump with $A_{0}=0.45$. Lower panel: rung-direction pump with $A_{0}=1$.
Figure 4: Connected charge (density-density) correlations $|\langle n_{i,l}n_{i+x,l}\rangle_{c}|$ along a leg of an infinite ladder as a function of separation $x$ and time $t$. The white dashed line indicates the time at which the pump is turned off. Upper panel: leg-direction pump with $A_{0}=1$. Middle panel: leg-direction pump with $A_{0}=0.45$. Lower panel: rung-direction pump with $A_{0}=1$.
@@
The charge correlation length in the ground state of the half filled Hubbard ladder with $U=8$ is much shorter than the spin correlation length, and this is evident in Fig. 4 at $t=0$. In contrast to the spin correlations, the spatial charge correlations only grow with time during the pump. This growth is strongest for the $A_{0}=1$ leg-direction pump (top panel), and weakest for the pump along the rung-direction. The growth proceeds without any apparent sharp features associated with the time at which the pump amplitude peaks or the time at which the pump ends. Unlike the spin correlations, the charge correlations within the ‘light cone’ appear to decrease monotonically for small $x$, and there is no distinct minimum at $x=2$ in the charge correlations for the less energetic pump protocols. Hence, there are distinct qualitative differences between the spatial spin and charge correlations, despite the simple relation between the local magnetisation and doublon number.
@@
Table 1: Correlation lengths for spin, $\xi_{s}$, and charge, $\xi_{c}$, in equilibrium and at $t=10$ after pumping. We fit the correlation length between $x=6$ and $x=12$ using the data shown in Fig. 5. We include 2 significant figures, fit uncertainties are $10^{-2}$.
@@
The spin correlation length $\xi_{s}$ is reduced by approximately one lattice unit after the pump. However, the most notable difference between the equilibrium and nonequilibrium spin correlations is the very fast decay of the latter at short range, before the exponential decay characterised by $\xi_{s}$ takes over. This behaviour is evident for $x\leq 3$ in the upper panel of Fig. 5 where, although the spin correlations at $x=0$ (i.e. the squared magnetizations) are of similar magnitude for the equilibrium and nonequilibrium data, by $x=3$ there is an order of magnitude difference between the equilibrium and nonequilibrium cases, showing strong suppression of magnetic correlations by the pump. This effect is even more pronounced for the $A_{0}=0.45$ leg-direction pump and the rung-direction pump because of the local minimum at $x=2$. Thus, the value of $\xi_{s}$ alone does not capture the extreme weakness of correlations beyond nearest neighbour in the nonequilibrium state.
The inset of the lower panel of Fig. 5 and the data in Table 1 show that connected charge correlations after pumping are strongly enhanced relative to equilibrium. Although the post pump spin and charge correlation lengths are similar, their short range behaviours are different. For charge correlations there is no additional strong suppression on going from $x=1$ to $x=2$, and there is no local minimum at $x=2$. The increase in $\xi_{c}$ can be interpreted as a consequence of the conservation of charge: the pump increases the doublon occupancy of a site, which must also reduce the surrounding charge density, and thus leads to an enhanced anticorrelation.
@@
In Fig. 6 we show the correlations at the antiferromagnetic wave vector $(\pi,\pi)$. We also include data for finite, $L=16$, rung ladders with open boundary conditions, using the two central rungs as a proxy for the two rung unit cell of approximate infinite system. The leg-direction $A_{0}=1$ pump shows the fastest suppression of the antiferromagnetic peak, with $S^{zz}(\pi,\pi,t)$ essentially reaching its late time value by $t\approx 2$, during the pump. For the lower fluence, $A_{0}=0.45$, leg-direction pump the suppression of $S^{zz}(\pi,\pi,t)$ is slower, reaching its late time value close to $t=\pi$ when the pump ends. The rung-direction pump leads to the slowest suppression of the three protocols, with $S^{zz}(\pi,\pi,t)$ decreasing until $t\approx 4$ after the pump has ended. At $t=0$ the $L=16$ rung data differ from the $L=\infty$ value of $S^{zz}(\pi,\pi,t)$ due to finite size effects. As the spin correlation length decreases, the difference between a finite and infinite system becomes less distinct, and for all three pump protocols there is good agreement between the $L=16$ and $L=\infty$ data by time $t\approx 1$. This agreement continues to improve until $t\approx 5$, at which point small differences between the $L=16$ and $L=\infty$ curves become apparent again. This is consistent with the partial recovery of longer range staggered spin correlations seen in Fig 3 which increases the influence of the boundary in the finite system. We note that there is still very good agreement at $t=10$ between the $L=16$ and $L=\infty$ data for the three protocols. As $\xi_{c}\approx\xi_{s}$, these results suggest that an $L=16$ system is a good approximation for an infinite ladder during and after the pump.
@@ IV Nonequilibrium dynamical structure factor
Here the $g(t_{i},t)$ are windowing functions associated with the probe pulse, which we assume to have a Gaussian form,
@@
where $\sigma_{X}$ is the probe’s width in time. The finite duration of the probe pulse limits the resolution of the time resolved response. $C(\mathbf{k},t_{1},t_{2})=\langle O_{\mathbf{k}}(t_{1})O^{\dagger}_{-\mathbf{k}}(t_{2})\rangle$ is a two time correlation function in momentum space. For our purposes $O_{\mathbf{k}}=S^{z}_{\mathbf{k}}$ or $n_{\mathbf{k}}$, respectively the spin $z$-component or charge density operator in momentum space.
@@
$\times\left[C(\mathbf{k},t+\tau,t)+C(\mathbf{k},t,t-\tau)\right].$ (6)
This approximation is reasonable if $C(\mathbf{k},t+\tau,t)$ varies slowly with $t$ relative to the scale $\sigma_{X}$. In our results we set $\sigma_{X}=1$. We provide more details in Appendix D including a comparison that shows there is good agreement between the approximate form Eq. 6 and the full form Eq. 4, even during the pump when the conditions for the approximation are not strictly met.
@@ IV.1 Nonequilibrium sum rules
In this subsection we show that the combined total spectral weight of the nonequilibrium dynamical spin and charge responses is time-independent. Although we concentrate on the case of the two leg ladder, the derivation generalizes in a simple manner and the result also applies to single band Hubbard models on other lattices (and in other dimensions). Relations of this type, between sums of dynamical quantities and a constant, are called sum rules.
The total spectral weight in a channel is given by integrating Eq. 4 for $I(\mathbf{k},\omega,t)$ over all $\mathbf{k}$ and $\omega$,
@@ IV.2 Spin dynamical structure factor
Figure 7: Nonequilibrium dynamical spin structure factor at $k_{y}=\pi$ for a pump with $A_{0}=1$ in the leg ($\parallel$) direction. The leftmost column shows the response in equilibrium before the pump. The middle and rightmost columns show results during and after the pump respectively. The black open circles mark maxima in the response at a given wavevector $k_{x}$. The corresponding plots for the leg-direction pump with $A_{0}=0.45$ and rung-direction pump with $A_{0}=1$ are qualitatively very similar.
We compute the dynamical spin response $S(k_{x},k_{y},\omega,t)$ using Eq. 6 before pumping, mid way through the pump at $t=\pi/2$ and after the pump at $t=10$.
@@
In equilibrium, the spin excitations of the Hubbard ladder at half filling and large $U$ are well described by those of a Heisenberg ladder with antiferromagnetic exchange interactions. This two-leg spin ladder has a ground state characterised by singlet formation on rungs [42]. Spin excitations are formed by breaking rung singlets, leading to massive (gapped) dispersing triplets. There is a spin gap and no true long range order, and the system is sometimes described as a spin liquid.
Parity selection rules mean that the spin response at different $k_{y}$ is related to the number of rung triplet excitations that are formed [55]. A rung singlet has negative parity (the wavefunction changes sign under interchange of the legs) whilst a rung triplet has positive parity (the wavefunction sign remains the same under interchange of the legs). If the ladder is in a state of well defined total parity then breaking an odd number of singlets to form triplets will change the total parity of the state from positive to negative or vice-versa. In contrast, breaking an even number of singlets to form triplets leaves the total parity unchanged. The spin operator $S^{z}_{k_{x},k_{y}}$ with $k_{y}=0$ preserves parity but for $k_{y}=\pi$ it alters parity. Accordingly, the spin response for $k_{y}=\pi$ probes the formation of odd numbers of triplets (as the parity is altered) and for $k_{y}=0$ it probes the formation of even numbers of rung triplets (as the total parity is unchanged). As the triplets are massive the minimum energy spin response is expected for $k_{y}=\pi$ which allows for the creation of a single rung triplet. However, because the spin gap $\Delta_{s}\approx 0.12t_{\parallel}$ is small and the antiferromagnetic correlations are somewhat long ranged, the low energy dynamics are approximately those of a spin-1 magnon, with a sharply defined mode that has an almost linear dispersion relation as it approaches zero energy near $\mathbf{k}=(\pi,\pi)$ [56, 57].
Fig. 7 shows the spin response for $k_{y}=\pi$ as a function of $\omega$ and $k_{x}$ for a leg-direction pump with $A_{0}=1$ at three different times: t=10t=-10 before pumping (equivalent to the equilibrium result), during the pump at $t=\pi/2$, and at $t=10$ after pumping. We do not show the results for the $A_{0}=0.45,\parallel$ and $A_{0}=1,\perp$ pumps as they are qualitatively very similar when plotted in this way. The equilibrium behaviour (convolved with the resolution due to the probe pulse’s temporal width) is seen in the left-hand panel of Fig. 7. The middle panel of the same figure shows that during the pump the spin response near $\mathbf{k}=(\pi,\pi)$ is reduced in intensity, in agreement with the real time results of the previous section. There is also a notable broadening of the response. At later times after pumping (right-hand panel of Fig. 7) the antiferromagnetic peak is substantially reduced in intensity, the response is broader for all $k_{x}$, and the dispersion has flattened.
@@
Figure 8: Nonequilibrium dynamical spin structure factor at $k_{y}=0$ for the $A_{0}=1$ leg-direction pump, before, during, and after the pump. The black open circles mark maxima in the response at a given wavevector $k_{x}$. The corresponding plots for the leg-direction pump with $A_{0}=0.45$ and rung-direction pump with $A_{0}=1$ are qualitatively very similar.
@@
Fig. 8 shows the equivalent dynamical spin response for the $A_{0}=1$ leg-direction pump along $k_{y}=0$ ( the other two pump protocols give qualitatively similar results). The response along $k_{y}=0$ vanishes as $k_{x}$ approaches 0 (because $\sum_{i,l}S^{z}_{i,l}=0$ is a conserved quantity) and is strongest at $k_{x}=\pi$. After the pump the response at $(\pi,0)$ is weaker, although this suppression is less significant than at $(\pi,\pi)$. There is also some enhancement of the spectral weight close to $k_{x}=0$ after pumping, relative to the equilibrium case.
@@
Figure 9: Magnetic correlations with $\mathbf{k}=(\pi,\pi)$, at $t=10$ after pumping along the legs $(\parallel)$ or rungs $(\perp)$. The inset shows the strong antiferromagnetic peak in equilibrium.
@@
Figure 10: Magnetic correlations with $\mathbf{k}=(\pi,0)$, at $t=10$ after pumping along the legs $(\parallel)$ or rungs $(\perp)$.
@@
More precise information on changes in the energy profile of the spin response is provided by taking cuts at fixed wave vectors. In Figs. 9 and 10 we compare the equilibrium and $t=10$ post pump responses for $\mathbf{k}=(\pi,\pi)$ and $\mathbf{k}=(\pi,0)$ respectively  [^1] The very small amount of negative spectral weight visible for two of the curves in Figs. 9 and 10 is a consequence of the finite window used in the frequency Fourier transform. This negative weight could be removed by using a larger time window at a significant computational cost.. The equilibrium response at both wave vectors is essentially a Gaussian peak, i.e. a single mode that is resolution limited due to the windowing effect of the probe pulse. The equilibrium response for $k_{y}=\pi$ can be interpreted in the rung triplet picture as showing interband transitions from the ground state to a band of single triplet excitations (or a band of double triplet excitations for $k_{y}=0$).
The most notable result of the pump at $\mathbf{k}=(\pi,\pi)$, other than the overall weakening of the response, is that the broadening of the peak occurs in an asymmetric manner with a tail towards higher energies, in contrast to the symmetric broadening of a single particle mode that is expected in a simple scattering lifetime approximation.
@@
Asymmetric broadening of a single particle mode in the dynamical spin response has been described previously in gapped spin systems with triplet excitations and hard-core interactions, at low but finite temperatures [59, 31, 32, 33, 60]. At zero temperature the single particle mode in these systems is a result of interband transitions from the ground state i.e. breaking a singlet and creating a state with a single propagating triplet. As temperature increases, triplet states become thermally populated and due to strong interactions they dress the single particle mode in a nontrivial manner, leading to an interband response with an asymmetric lineshape. Broadening of the single particle response occurs due to interband transitions from a thermally populated single particle state to a state with two particles. The precise form of the joint density of states for occupied single particle modes and unoccupied two particle states determines the direction of the asymmetry. Moreover, the existence of thermally populated triplets reduces the joint density of states for interband transitions, so the spectral weight associated with the single mode response decreases. These effects are accompanied by the emergence of a new spin response at $\omega\sim 0$ due to intraband transitions (scattering of the probe from thermally occupied states) which can occur for vanishing values of energy and momentum. This increased spectral weight at low energies is known as a Villain mode [61], and it compensates for the decrease in the single mode weight. The total integrated spin response then remains a conserved quantity as expected for a pure spin model.
@@
The total spin spectral weight $W_{S^{z}}$ can be determined using Eq. 10 from the doublon number computed with iTEBD, or by integrating the dynamical spin response for the 16 rung system shown in Figs. 7 and  8 over all $\omega$ and $\mathbf{k}$. We find excellent agreement between both these methods, which provides further confirmation of the accuracy of the approximation Eq. 6 and the use of the 16 rung system. For the $A_{0}=1$ leg-direction pump the ratio of the spin weight after the pump to that in equilibrium is $W_{S^{z}}(t=10)/W_{S^{z}}(t=-10)=0.76$. For the $A_{0}=1$ rung-direction pump the corresponding figure is $W_{S^{z}}(t=10)/W_{S^{z}}(t=-10)=0.86$. Therefore, the integrated spin response decreases significantly after the optical pump, indicating a substantial transfer of spectral weight to the charge sector.